OpenStudy (anonymous):

What is the solution of the sollowing system of equations? 3x+2y+z=7 5x+5y+4z=3 3x+2y+3z=1

OpenStudy (anonymous):

would i multiply the first equation by -1 and then add that and the third? -3x-2y-x=-7 3x+2y+3z=1 ?

OpenStudy (mosaic):

Yes. That would get rid of the x variable using equations 1 and 3. Eliminate the x variable again but this time using equations 1 and 2. Multiply the first equation by 5 and the second equation by -3 and add them to get rid of x. You will now have two equations and two unknowns y and z. Eliminate one of the variables and solve.

OpenStudy (anonymous):

yeah do it..

OpenStudy (anonymous):

but that would eliminate the x and the y?

OpenStudy (mosaic):

True which makes life a lot easier! You can straightaway solve for z!

OpenStudy (mosaic):

3x+2y+z=7 (1) 5x+5y+4z=3 (2) 3x+2y+3z=1 (3) Multiply (1) by -1 and add to (3): -3x-2y-z=-7 3x+2y+3z=1 add 2z = -6 z = -3 Put z = -3 in (1) and (2) and solve for x and y.

OpenStudy (anonymous):

im a little confused, i put z into either of the equations to solve for x and the for y?

OpenStudy (mosaic):

Put z = -3 in (1): 3x+2y+(-3)=7 3x+2y=10 (4) Put z = -3 in (2): 5x+5y+4(-3)=3 5x+5y-12=3 5x+5y=15 x + y = 3 (5) Use equations (4) and (5) to solve for x and y. It is two simultaneous equations and two unknowns.

OpenStudy (anonymous):

that just confused me more,

OpenStudy (mosaic):

We started with three equations and three unknowns. Then we reduce it to two equations and two unknowns: 3x+2y=10 (4) x + y = 3 (5) multiply equation (5) by -2 and add it to equation (4): 3x+2y=10 -2x-2y=-6 add x = 4 Put it in (5) and solve for y: 4+y=3 y = -1 x = 4, y = -1, z = -3.